A well-known French fashion house and luxury retail company, Louisa Vuittona founded in 1855 produces luggage and handbags as its signature pieces, which are handmade in France. For their handbag production in western France, they have, on average, 44,000 handbags in production. They ship 12,000 bags on average every week (52 weeks a year) to retail outlets around the world. There are no finished goods left in the factory after each daily shipment. What is the average length of time (in weeks) a handbag takes to be produced in the Luisa Vuittona factory?

Select one:

a. 0.05

b. 0.89

c. 11.0

d. 18.6

e. 20.9
Answer #1

Correct Answer: 3.67 weeks

Correct answer is not given in the options

WIP inventory, W = 44,000 bags

Throughput, R = 12000 bags per week

As per Little's law, W = R*T

So, average length of time to produce, T = W/R  

= 44000/12000

= 3.67 weeks